A truck and plane leave the same area at the same time and head for a town 450 miles away. The plane is three times the speed of the truck and arrives 6 hours ahead of the truck. Find the speed of the truck.
Answer by mananth(16949)

A truck and plane leave the same area at the same time and head for a town 450 miles away. The plane is three times the speed of the truck and arrives 6 hours ahead of the truck. Find the speed of the truck.
..
let speed of truck be xmph
speed of plane = 3x
..
Distance traveled by both = 450 miles.
..
Time taken by truck = 450/x
Time taken by plane = 450/3x
..
Time taken by truck - time taken by plane = 6
450/x - 450/3x= 6
LCM = 3x
1350-450 /3x= 6
900 / 3x= 6
18x=900
x=900/18
x=50 mph --- the speed of the truck
CHECK
450/50- 450/150
9-3=6 hours